OpenStudy (anonymous):

.Find the distance between the given points: (4√15, -3√10) (2√15, 8√10)

sam (.sam.):

Given the two points \((x_1, y_1)\) and \((x_2, y_2)\), the distance between these points is given by the formula: \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] Substitute and solve.

OpenStudy (anonymous):

im having trouble with the radicals could u solve this for?

sam (.sam.):

\[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2} \\ \\ d=\sqrt{(2\sqrt{15}-4\sqrt{15})^2+(8\sqrt{10}-(-3\sqrt{10}))^2} \\ \\ d=\sqrt{(-2\sqrt{15})^2+(11\sqrt{10})^2} \\ \\ d=\sqrt{(4(15))+(121(10))} \]

sam (.sam.):

Note: \(m\sqrt{x}+n\sqrt{x}=(m+n)\sqrt{x}\)

OpenStudy (anonymous):

so what would the answer be?

sam (.sam.):

Just finish up the solving part

OpenStudy (anonymous):

ok thanks

sam (.sam.):

yw